The length of a side of an equilateral triangle is the same as the length of a rectangle and the width of the rectangle is 2 inches less than its length. If the perimeter of the triangle is 4 inches less than the perimeter of the rectangle, what are the dimensions of the rectangle?
Answer by jorel1380(3719)

2(l)+2(l-2)=3l
2l+2l-4=3l
4l-4=3l
l=4
w=2
The rectangle is 4 inches by 2 inches..